OpenStudy (anonymous):

Given the function f(x) = −2x + 8, find x if f(x) = 22. −36 −7 52 −15

jimthompson5910 (jim_thompson5910):

f(x) = −2x + 8 22 = −2x + 8 22-8 = −2x 14 = −2x 14/(-2) = x -7 = x x = -7